OpenStudy (anonymous):

sin theta 3 /5 find the value of tan θ

OpenStudy (eric_d):

tan theta = sin theta/ cos theta

OpenStudy (neer2890):

\[\sin \theta=\frac{ Perpendicular }{ Hypotenuse }\]  using pythagorean theorem in above triangle \[x ^{2}=5^{2}-3^{2}\] \[x ^{2}=16\] x=4 now, \[\tan \theta=\frac{ Perpendicular }{ Base }\] \[\tan \theta=\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

\[\tan \left(\sin ^{-1}\left(\frac{3}{5}\right)\right)=\frac{3}{4} \]